OpenStudy (anonymous):

Which equation is an identity? a. 4m-4=5m+5-m b. 14-(4v+7)=-4v - 7 c. 5y+3=5y-3 d. 5w+6-w=6w-2(w-3)

OpenStudy (king):

option d

OpenStudy (radar):

all of em

OpenStudy (radar):

seems like they all end up with the variable disappearing!

jhonyy9 (jhonyy9):

d.

OpenStudy (radar):

a.4m-4=5m+5-m 4m-5m+m=5+4 0m=9 b.14-4v-7=-4v-7 -4v+4v=-7-14+7 0v=-14 c.5y+3=5y-3 5y-5y=-3-3 0y=-6 d.5w+6-w=6w-2w+6 5w-w-6w+2w=6-6 0w=0 All of the equations are unsolvable because division by 0 is undefined. Why pick on d. as the bad boy? Is it because the left and right is identical in the beginning?